number of atoms in 560g of fe(atomic mass 56g mol inverse)is:

Grade:12th pass

1 Answers

Ritik
6 years ago
Number of mole of fe =560/56=10. Number of atom in one mole=6.022× 10^23 So number of atom in 10 mole of fe = 60.22×10^23